solution for the area of rectangle with sides b yards and b - 6 yards
Answer by stanbon(75887)

solution for the area of rectangle with sides b yards and b - 6 yards
---
Area = b(b-6) = b^2-6b sq. yds.
====================================
Cheers,
Stan H.